OpenStudy (anonymous):

y varies directly as the square of x. When x = 5, y = 5. Find y when x = 10.

OpenStudy (anonymous):

is y=(1/5)*x*x if x=10 y=20

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

use y= kx2 and after putting for x=5 and y=5 we have k=1/5

OpenStudy (callisto):

y α x^2 y/x^2 = k \[\frac{y_1}{x_1\ ^2} = \frac{y_2}{x_2\ ^2}\]Put \(x_1=5, y_1=5\) and \(x_2 = 10\)into it. \[\frac{5}{5^2} = \frac{y}{10^2}\]\[\frac{1}{5} = \frac{y}{100}\]Solve y.

OpenStudy (anonymous):

oh wow that makes alot of sense thanks

OpenStudy (callisto):

Welcome!